OpenStudy (anonymous):

which ratio is proportional to 8/18? 2/5 4/5 16/26 20/45

OpenStudy (anonymous):

@563blackghost @pooja195

OpenStudy (anonymous):

i have no idea how to do this......

OpenStudy (anonymous):

@jabez177

563blackghost (563blackghost):

When two quantities always have the same size in relation to each other. In other words they have the same ratio. cited~ https://www.mathsisfun.com/definitions/proportional.html

OpenStudy (danjs):

it just wants to know which ones are the same as that given one, right?

OpenStudy (anonymous):

Ummmm..............I think so...............

OpenStudy (danjs):

some constant times your fraction \[k*\frac{ 8 }{ 18 }\]

OpenStudy (anonymous):

I think it might be A

OpenStudy (danjs):

they gave you answers to pick set them each equal to the given one, and cross multiply and see if it is true

563blackghost (563blackghost):

Incorrect.... we first would find the simplified fraction of the fraction in the question.... so factors of 8 are 1,2,4,8 while factors of 16 are 1, 2, 3, 6, 9, 18

OpenStudy (danjs):

\[\frac{ 2 }{ 5 }=\frac{ 8 }{ 18 }\] 40 = 36, nope, not one of the answers

OpenStudy (anonymous):

Ohhh ok

563blackghost (563blackghost):

Now we choose on in which both have the same highest factor....so which factor that is the largest do they both contain...

OpenStudy (anonymous):

ekkkk................IDK

OpenStudy (anonymous):

8 and 18 are the highest

OpenStudy (danjs):

does 20/45 = 8/18 cross multiply 20*18 = 45*8 360 = 360 true

OpenStudy (anonymous):

wait is the answer C

OpenStudy (anonymous):

ohhhhh

563blackghost (563blackghost):

They both are divisible by 2 so \[\frac{ 8 }{ 18 }\div \frac{ 2 }{ 2 }=\frac{ 4 }{ 9 }\]

jabez177 (jabez177):

^ Exactly what I was going to do. :)

OpenStudy (anonymous):

wait so..........Im sorta confuesed

563blackghost (563blackghost):

So we are able to count out A and B since they are both already in their simplified forms....so all we need to do is simplify C and D....

jabez177 (jabez177):

Just divide the numerator and the denominator by 2 and you will get your answer. :)

jabez177 (jabez177):

Wait. Nvm.

OpenStudy (danjs):

i would just set 8/18 = to each choice, see if its true by cross multiplying,

OpenStudy (danjs):

or 4/9 if you will

OpenStudy (anonymous):

idk what to do???????

OpenStudy (anonymous):

@Jaynator495

563blackghost (563blackghost):

sorry my internet is not working right...one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

yo ^-^

OpenStudy (just_one_last_goodbye):

is it the first question?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

in the same proportion, proportional fractions here, means which are the same

OpenStudy (danjs):

in the same ratio

OpenStudy (danjs):

\[\frac{ a }{ b }=\frac{ c }{ d }\] does a*d = b*c if yes, they are proportional fractions

OpenStudy (anonymous):

so 16/26 doesn't wokr

OpenStudy (anonymous):

*work

OpenStudy (tkhunny):

Reduce to lowest terms and just look at them. 8/18 = 4/9 (reduced by 2) 2/5 = 2/5 (already minimal) 4/5 = 4/5 (already minimal) 16/26 = 8/13 (reduced by 2) 20/45 = 4/9 (reduced by 5) Find common factors and reduce until there are none.

OpenStudy (anonymous):

it isn't 16/26 it is 16/36 so yea if you doing 8/18

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

I am telling the truth none of those things you asked are right

OpenStudy (anonymous):

I'm still connfuesed

OpenStudy (anonymous):

I am telling you this 18/36 would be the right answer

OpenStudy (anonymous):

you got a proportion for example 1/2 = 2/4

OpenStudy (anonymous):

but that's not the answer chioce

OpenStudy (anonymous):

that is an example

OpenStudy (anonymous):

the answer is 18/36

OpenStudy (anonymous):

@malcolmmcswain

OpenStudy (anonymous):

or 4/9 ask teacher

OpenStudy (anonymous):

or a multiple of 16/36

OpenStudy (anonymous):

I am teaching my teacher in class all the time

OpenStudy (anonymous):

@tygrr321

OpenStudy (anonymous):

my teacher gave me a 73% and I need to start telling her the right answers and what not

OpenStudy (anonymous):

she ain't smart

OpenStudy (anonymous):

I'm going with D

OpenStudy (anonymous):

as I thought with checking

OpenStudy (anonymous):

well

OpenStudy (tkhunny):

@sydneymmelin Stop tagging people and start showing your work. Reduce this fraction. 2/4 If you cannot do it, you cannot solve this problem, except by throwing darts and getting lucky.

OpenStudy (anonymous):

I guess so

OpenStudy (anonymous):

that would be correct but not really right in a sense

OpenStudy (anonymous):

D is correct though in a sense

OpenStudy (anonymous):

not in all ways of teaching it is correct depends on what your teacher says

OpenStudy (anonymous):

but go for it

OpenStudy (anonymous):

I am in middle school so I don't have too much work

OpenStudy (anonymous):

I am going with 20/45 bc if you cross multiply 8/18 like @DanJS said you would get 360 on both sides

OpenStudy (anonymous):

true

OpenStudy (anonymous):

I guess you go with that

OpenStudy (anonymous):

it isn't right but do it since it is the closest to correct answer

OpenStudy (anonymous):

since it is usually a multiple or a factor of the main fraction but it is still correct

OpenStudy (anonymous):

can I get medal?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

because you didn't help much

OpenStudy (tkhunny):

That's okay, Liam. Sydney won't get a medal either, because even less help came from the OP. Next time, Sydney, show YOUR work.

OpenStudy (anonymous):

LOL I gave dan a medal bc he helped

OpenStudy (tkhunny):

Except for tagging people, this is essentially your contribution to this discussion: I have no idea how to do this...... I think it might be A ekkkk................IDK 8 and 18 are the highest wait is the answer C wait so..........Im sorta confused idk what to do??????? so 16/26 doesn't work I'm still connfuesed but that's not the answer choice I'm going with D I am going with 20/45 bc if you cross multiply 8/18 like @DanJS said you would get 360 on both sides This is a standard pattern for getting other people to do your work for you. Do better.

OpenStudy (danjs):

yeah, not as bad as,, "I think it is (Letter)" .. hah..

OpenStudy (anonymous):

wow thnkas

OpenStudy (anonymous):

thanks someones not getting a nedal

OpenStudy (tkhunny):

Just the facts. I have no need to make any other presentation.